OpenStudy (anonymous):

convert the rate 1000 ounces per 2 hours to cups per minutes?

OpenStudy (wolf1728):

8 ounces = 1 cup 1000 ounces = (1000/8) cups or 125 cups 2 hours = 2 * 60 = 120 minutes divide the 125 cups by 120 and so we have 1.0416666667 cups per minute